TIME RESOLVED INFRARED DIODE LASER SPECTROSCOPY OF THE ν_{1} BAND OF CoNO PRO-DUCED BY THE ULTRAVIOLET PHOTOLYSIS OF Co(CO)_{3}NO

\underline{TAKEO SOEJIMA}, MOTOKI NAKASHIMA, SEIKI IKEDA and KEIICHI TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

Rovibrational transitions of the ν_{1} band (N-O stretch) of CoNO produced by the ultraviolet photolysis of Co(CO)_{3}NO were observed in the 1775-1800 cm^{−1} region by time resolved infrared diode laser spectroscopy. In total, 35 absorption lines were assigned to the ν_{1} fundamental band, and their J quantum numbers were determined by a simultaneous analysis with millimeter-wave spectroscopy^{a}. Hotband lines originated from the ν_{2} (bending), 2ν_{2}, and ν_{3} (Co-N stretch) vibrationally excited states were also observed. Molecular constants, including the band origin ν_{0} 1796.22371(49) cm^{−1}, the rotational constant B_{0} 4669.7578(29) MHz, and the vibration rotation constant α_{1} 31.325(28) MHz, were derived from the observed spectrum. The equilibrium rotational constant B_{e} (4676.949(51) MHz) was determined with the α_{1} value derived in the present study and the α_{2} and α_{3} values reported by the millimeter-wave spectroscopy^{a}. The equilibrium bond-length r_{Co−N} was calculated to be 1.583 Å assuming r_{NO} = 1.182 Å by ab initio calculation^{b}. The absorption lines for the ν_{1} + ν_{2} ← ν_{2} and ν_{1} + 2ν_{2} ← 2ν_{2} bands were split into two components due to the Δl=2 interaction. The electronic ground state of CoNO was confirmed to be ^{1}Σ^{+} by this infrared study as reported by our previous millimeter-wave spectroscopy^{a}.

aTanaka et al, the 59th OSU International Symposium on Molecular Spectroscopy (2004).

bM. Zhou and L. Andrews, J. P hys. Chem. A 104, 3915 (2000).